# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 1.4. Organization of the article
The organization of this article is as follows: In the next section, we provide necessary function spaces, some important inequalities, linear and nonlinear operators with their properties for abstract formulation of system (1.5) and further analysis and abstract formulation of system (1.5). The metric dynamical system (MDS) and RDS corresponding to stochastic third-grade fluids equations are constructed in Section 3 using a random transformation (known as Doss-Sussman transformation, [29, 59]). The existence and uniqueness of a weak solution satisfying the energy equality to the transformed stochastic third-grade fluids equations (see system (3.19) below) by using monotonicity arguments and a Minty-Browder technique has also been established (Theorem 3.8) in the same section. Section 4 provides the existence of a unique random attractor for stochastic third-grade fluids equations on bounded domains via compact Sobolev embeddings and abstract theory developed in [13] (see Lemma 4.2 and Theorem 4.3). In Section 5, we consider the stochastic third-grade fluids equations on unbounded domains. In order to do this, we first present Lemmas 5.1 and 5.2, which provides us the energy estimates for the solution of system (3.19) and helps us to prove the existence of random absorbing set. Then, we present Lemma 5.3 which is one of the key lemma to prove the $\mathfrak{DK}$-asymptotic compactness of RDS. Next we prove the uniform-tail estimates for the solution of system (3.19) which is another key result to prove the $\mathfrak{DK}$-asymptotic compactness of RDS (see Lemma 5.4). Finally, we establish the main result of this section, that is, the existence of unique random attractor for system (1.5) on unbounded domains using Lemmas 5.2, 5.3 and 5.4, and abstract theory developed in [13], see Theorem 5.5. In the final and smallest section, we address the existence of invariant measures for system (1.5) as a consequence of the existence of random attractor (see Theorem 6.2)
@@ 2.1. Notations and the functional setting
$\begin{array}[]{ll}\mathscr{V}&:=\{\boldsymbol{v}\in(\mathrm{C}^{\infty}_{c}(\mathscr{O}))^{d}\,:\text{ div }\boldsymbol{v}=0\},\\[2.84526pt] \mathbb{H}&:=\text{The closure of }\mathscr{V}\text{ in }\mathbb{L}^{2}(\mathscr{O})=\{\boldsymbol{v}\in\mathbb{L}^{2}(\mathscr{O})\,:\text{ div }\boldsymbol{v}=0\text{ in }\mathscr{O}\;\;\text{and}\;\;\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial\mathscr{O}\},\\[2.84526pt] \mathbb{V}&:=\text{The closure of }\mathscr{V}\text{ in }\mathbb{H}^{1}(\mathscr{O})=\{\boldsymbol{v}\in\mathbb{H}^{1}_{0}(\mathscr{O})\,:\text{ div }\boldsymbol{v}=0\text{ in }\mathscr{O}\}\\ \mathbb{V}_{s}&:=\text{the closure of }\ \mathscr{V}\ \text{ in the Sobolev space }\mathbb{H}^{s}(\mathscr{O}),\ \text{ for }s>1.\end{array}$
Now, we recall the Leray-Helmholtz projector $\mathscr{P}:\mathbb{L}^{2}(\mathscr{O})\to\mathbb{H}$, which is a linear bounded operator characterized by the following $\mathbb{L}^{2}$-orthogonal decomposition $\boldsymbol{v}=\mathscr{P}\boldsymbol{v}+\nabla\varphi,\;\varphi\in\mathrm{H}^{1}(\mathscr{O})$, (cf. [61]).
Now, let us introduce the scalar product between two matrices $A:B=\mathrm{Tr}(AB^{T})$ and denote $|A|^{2}:=A:A.$ The divergence of a matrix $A\in\mathscr{M}_{d\times d}(E)$ is given by $(\text{div}(A)_{i})_{i=1}^{i=d}=(\displaystyle\sum_{j=1}^{d}\partial_{j}a_{ij})_{i=1}^{i=d}.$ The space $\mathbb{H}$ is endowed with the $\mathbb{L}^{2}$-inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and the associated norm $\|\cdot\|_{2}$. We recall that
@@
$(A,B)$ $=\int_{\mathscr{O}}A:B\mathrm{d}x;\quad\text{ for all }A,B\in\mathscr{M}_{d\times d}(L^{2}(\mathscr{O})).$
On the functional space $\mathbb{V}$, we will consider the following inner product
$\begin{array}[]{ll}(\boldsymbol{u},\boldsymbol{v})_{\mathbb{V}}&:=(\boldsymbol{u},\boldsymbol{v})+(\nabla\boldsymbol{u},\nabla\boldsymbol{v}),\end{array}$
and denote by $\|\cdot\|_{\mathbb{V}}$ the corresponding norm. The usual norms on the classical Lebesgue and Sobolev spaces $\mathrm{L}^{p}(\mathscr{O})$ (resp. $\mathbb{L}^{p}(\mathscr{O})$) and $\mathrm{W}^{m,p}(\mathscr{O})$ (resp. $\mathbb{W}^{m,p}(\mathscr{O})$) will be denoted by $\|\cdot\|_{p}$ and $\|\cdot\|_{\mathrm{W}^{m,p}}$ (resp. $\|\cdot\|_{\mathbb{W}^{m,p}}$), respectively. In addition, given a Banach space $E$, we will denote by $E^{\prime}$ its dual.
@@
[…]
Let us recall an embedding result from [44, Subsection 2.4].
@@
    Gagliardo-Nirenberg inequality ([51, Theorem 1]) gives (for $d=2,3$)
@@
    Korn-type inequality (see [28, Theorem 2 (ii)] and [46, Theorem 4, p.90]) gives that there exists a constant $C_{K}>0$ such that
@@ 2.2. Linear and nonlinear operators
$\mathscr{A}\boldsymbol{v}:=-\mathscr{P}\Delta\boldsymbol{v},\ \boldsymbol{v}\in\mathrm{D}(\mathscr{A}):=\mathbb{V}\cap\mathbb{H}^{2}(\mathscr{O}).$
@@
$\mathscr{A}\boldsymbol{v},\boldsymbol{v}\right>=\|\boldsymbol{v}\|_{\mathbb{V}}^{2},\ \textrm{ for all }\ \boldsymbol{v}\in\mathbb{V},\ \text{ so that }\ \|\mathscr{A}\boldsymbol{v}\|_{\mathbb{V}^{\prime}}\leq\|\boldsymbol{v}\|_{\mathbb{V}}.$ (2.3)
Next, we define the trilinear form $b(\cdot,\cdot,\cdot):\mathbb{V}\times\mathbb{V}\times\mathbb{V}\to\mathbb{R}$ by
$b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{w})=\int_{\mathscr{O}}(\boldsymbol{u}(x)\cdot\nabla)\boldsymbol{v}(x)\cdot\boldsymbol{w}(x)\mathrm{d}x=\sum_{i,j=1}^{d}\int_{\mathscr{O}}u_{i}(x)\frac{\partial v_{j}(x)}{\partial x_{i}}w_{j}(x)\mathrm{d}x.$
@@
{$b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{v})&=0,\ \text{ for all }\ \boldsymbol{u},\boldsymbol{v}\in\mathbb{V},\\ b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{w})&=-b(\boldsymbol{u},\boldsymbol{w},\boldsymbol{v}),\ \text{ for all }\ \boldsymbol{u},\boldsymbol{v},\boldsymbol{w}\in\mathbb{V}.\end{aligned}\right.$ (2.4)
The trilinear map $b:\mathbb{V}\times\mathbb{V}\times\mathbb{V}\to\mathbb{R}$ has a unique extension to a bounded trilinear map from $\mathbb{L}^{4}(\mathscr{O})\times\mathbb{W}^{1,4}(\mathscr{O})\times\mathbb{H}$ to $\mathbb{R}$. For any $\boldsymbol{u}\in\mathbb{W}^{1,4}(\mathscr{O})$ and $\boldsymbol{v}\in\mathbb{H}$, we have
@@
Let us now define the operator $\mathscr{J}(\boldsymbol{v}):=-\mathscr{P}\mathrm{div}(\mathrm{A}(\boldsymbol{v})\mathrm{A}(\boldsymbol{v}))$. Note that for $\boldsymbol{v}\in\mathbb{W}^{1,4}(\mathscr{O})$, we have
$\|\mathscr{J}(\boldsymbol{v})\|_{\mathbb{V}^{\prime}}\leq C\|\mathrm{A}(\boldsymbol{v})\|^{2}_{4},$
and hence the map $\mathscr{J}(\cdot):\mathbb{W}^{1,4}(\mathscr{O})\to\mathbb{V}^{\prime}$ is well-defined.
Finally, we define the operator $\mathscr{K}(\boldsymbol{v}):=-\mathscr{P}\mathrm{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))$. It is immediate that
@@
Furthermore, the operator $\mathscr{K}(\cdot):\mathbb{W}^{1,4}(\mathscr{O})\to\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ is well-defined.
@@
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$ and $\varepsilon_{0}:=1-\sqrt{\frac{\alpha^{2}}{2\beta\nu}}\in(0,1)$. Then, the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ given by
@@
$\mathscr{G}(\boldsymbol{y}_{1})-\mathscr{G}(\boldsymbol{y}_{2}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
@@
$\nu\mathscr{A}\boldsymbol{y}_{1}-\nu\mathscr{A}\boldsymbol{y}_{2},\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>=\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}.$ (2.7)
@@
|$\mathscr{B}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>|$ |
@@
$\mathscr{B}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>$
@@
From [34, Equation (2.21)], we have
$|\alpha\left<\mathscr{J}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>|$
@@
Now, from [34, Equation (2.13)], we have
@@
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$. Then, the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ is demicontinuous.
@@
Let us take a sequence $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{X}$, that is, $\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{X}}=\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{V}}+\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\to 0$ as $m\to\infty$. For any $\boldsymbol{\phi}\in\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\mathbb{X}}$, we consider
@@
$|\langle\mathscr{A}\boldsymbol{y}_{m}-\mathscr{A}\boldsymbol{y},\boldsymbol{\phi}\rangle|=|(\nabla(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi})|\leq\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{V}}\|\boldsymbol{\phi}\|_{\mathbb{V}}\to 0,\ \text{ as }\ m\to\infty,$ (2.10)
since $\boldsymbol{y}_{n}\to\boldsymbol{y}$ in $\mathbb{V}$. We estimate the term $\langle\mathscr{B}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.2) using (2.4), Hölder’s inequality and Lemma 2.1 as
@@
$\leq C\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C}\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}+\boldsymbol{z}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}$
$\to 0,\ \text{ as }\ m\to\infty,$ (2.11)
since $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{X}$. We estimate the term $\langle\mathscr{J}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.2) using Hölder’s inequality as
$|\alpha\langle\mathscr{J}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{J}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{y}+\boldsymbol{z}}),\boldsymbol{\phi}\rangle|$
@@
$\to 0,\ \text{ as }\ m\to\infty,$ (2.12)
since $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{W}^{1,4}(\mathscr{O})$. We estimate the term $\langle\mathscr{K}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.2) using Hölder’s inequality as
@@
$\to 0,\ \text{ as }\ m\to\infty,$ (2.13)
since $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{W}^{1,4}(\mathscr{O})$.
From the above convergences, it is immediate that $\langle\mathscr{G}(\boldsymbol{y}_{m})-\mathscr{G}(\boldsymbol{y}),\boldsymbol{\phi}\rangle\to 0$, for all $\boldsymbol{\phi}\in\mathbb{X}$. Hence the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ is demicontinuous, which implies that, in addition, the operator $\mathscr{G}(\cdot)$ is hemicontinuous also. ∎
@@
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$. Then, the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ is locally Lipschitz.

### v2

@@ 1.4. Organization of the article
The organization of this article is as follows: In the next section, we provide necessary function spaces, some important inequalities, linear and nonlinear operators with their properties for abstract formulation of system (1.5) and further analysis and abstract formulation of system (1.5). The metric dynamical system (MDS) and RDS corresponding to system (1.5) are constructed in Section 3 using a random transformation (known as Doss-Sussmann transformation, [30, 67]). The existence and uniqueness of a weak solution satisfying the energy equality to the transformed system (see system (3.19) below) by using monotonicity arguments and a Minty-Browder technique has also been established (Theorem 3.8) in the same section. Section 4 provides the existence of a unique random attractor for system (1.5) on bounded domains via compact Sobolev embeddings and abstract theory developed in [10] (see Lemma 4.2 and Theorem 4.3). In Section 5, we consider the stochastic system (1.5) on unbounded domains. In order to do this, we first present Lemmas 5.1 and 5.2, which provides us the energy estimates for the solution of system (3.19) and helps us to prove the existence of random absorbing set. Then, we present Lemma 5.3 which is one of the key lemma to prove the $\mathfrak{DK}$-asymptotic compactness of RDS. Next we prove the uniform-tail estimates for the solution of system (3.19) which is another key result to prove the $\mathfrak{DK}$-asymptotic compactness of RDS (see Lemma 5.4). Finally, we establish the main result of this section, that is, the existence of unique random attractor for system (1.5) on unbounded domains using Lemmas 5.2, 5.3 and 5.4, and abstract theory developed in [10], see Theorem 5.5. In the final and smallest section, we address the existence of invariant measures for system (1.5) as a consequence of the existence of random attractor (see Theorem 6.2). In Appendix A, we have shown that one can retrieve the pressure from the abstract formulation of system (1.5) using the approach discussed in [66, Section 8].
@@ 2.1. Notations and the functional setting
[…]
where $\boldsymbol{n}$ is the unit outward drawn normal to $\partial\mathscr{O}$, and $\boldsymbol{v}\cdot\boldsymbol{n}\big|_{\partial\mathscr{O}}$ should be understood in the sense of trace in $\mathbb{H}^{-1/2}(\partial\mathscr{O})$ (cf. Theorem 1.4 and Remark 1.5, Chapter 1, [70]).
Now, let us introduce the scalar product between two matrices $A:B=\mathrm{Tr}(AB^{T})$ and denote $|A|^{2}:=A:A.$ The divergence of a matrix $A\in\mathscr{M}_{d\times d}(E)$ is given by $(\text{div}(A)_{i})_{i=1}^{i=d}=(\displaystyle\sum_{j=1}^{d}\partial_{j}a_{ij})_{i=1}^{i=d}.$ The space $\mathbb{H}$ is endowed with the $\mathbb{L}^{2}$-inner product $(\cdot,\cdot)$ and the associated norm $\|\cdot\|_{2}$. We recall that
@@
[…]
In view of (1.8), on the functional space $\mathbb{V}$, we will consider the following inner product
$(\boldsymbol{u},\boldsymbol{v})_{\mathbb{V}}:=(\nabla\boldsymbol{u},\nabla\boldsymbol{v}),$
and denote by $\|\cdot\|_{\mathbb{V}}$ the corresponding norm. The usual norms on the classical Lebesgue and Sobolev spaces $\mathrm{L}^{p}(\mathscr{O})$ (resp. $\mathbb{L}^{p}(\mathscr{O})$) and $\mathrm{W}^{m,p}(\mathscr{O})$ (resp. $\mathbb{W}^{m,p}(\mathscr{O})$) will be denoted by $\|\cdot\|_{p}$ and $\|\cdot\|_{\mathrm{W}^{m,p}}$ (resp. $\|\cdot\|_{\mathbb{W}^{m,p}}$), respectively. In addition, given a Banach space $\mathrm{E}$, we will denote by $\mathrm{E}^{\prime}$ its dual, and for any Hilbert space $\mathrm{H}$, we represent inner product by $(\cdot,\cdot)_{\mathrm{H}}$. We denote by $\langle\cdot,\cdot\rangle$ the duality pairing between $\mathrm{E}^{\prime}$ and $\mathrm{E}$.
@@
Let us recall an embedding result from [48, Subsection 2.4].
@@
    Gagliardo-Nirenberg inequality ([57, Theorem 1]) gives (for $d=2,3$)
@@
    Korn-type inequality (see [29, Theorem 2 (ii)] and [51, Theorem 4, p.90]) gives that there exists a constant $C_{K}>0$ such that
@@
2.2. The Helmholtz projection
The content of this subsection has been inspired from the articles [33, 34, 52] which discuss the Helmholtz projection on general unbounded domains with $\mathrm{C}^{1}$-boundary. Let us define
$\widetilde{\mathbb{L}}^{p}(\mathscr{O}):=\begin{cases}\mathbb{L}^{p}(\mathscr{O})\cap\mathbb{L}^{2}(\mathscr{O}),&p\in[2,\infty)\\ \mathbb{L}^{p}(\mathscr{O})+\mathbb{L}^{2}(\mathscr{O}),&p\in(1,2),\end{cases}$ (2.3)
and
$\widetilde{\mathbb{L}}^{p}_{\sigma}:=\begin{cases}\mathbb{L}^{p}_{\sigma}\cap\mathbb{L}^{2}_{\sigma},&p\in[2,\infty)\\ \mathbb{L}^{p}_{\sigma}+\mathbb{L}^{2}_{\sigma},&p\in(1,2).\end{cases}$ (2.4)
Now, we recall that the Helmholtz projector $\mathscr{P}_{p}:\widetilde{\mathbb{L}}^{p}(\mathscr{O})\rightarrow\widetilde{\mathbb{L}}^{p}_{\sigma}$, which is a linear bounded operator characterized by the following decomposition $\boldsymbol{v}=\mathscr{P}_{p}\boldsymbol{v}+\nabla\varphi,\;\varphi\in\widetilde{\mathrm{G}}^{p}(\mathscr{O})$ satisfying $(\mathscr{P}_{p})^{*}=\mathscr{P}_{p^{\prime}}$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$ (cf. [34]), where
$\widetilde{\mathrm{G}}^{p}(\mathscr{O}):=\begin{cases}\mathrm{G}^{p}(\mathscr{O})\cap\mathrm{G}^{2}(\mathscr{O}),&p\in[2,\infty)\\ \mathrm{G}^{p}(\mathscr{O})+\mathrm{G}^{2}(\mathscr{O}),&p\in(1,2),\end{cases}$ (2.5)
and
$\mathrm{G}^{p}(\mathscr{O})=\{\nabla q\in\mathbb{L}^{p}(\mathscr{O})\;:\;q\in\mathrm{L}^{p}_{loc}(\mathscr{O})\}.$ (2.6)
Let us fix the notations $\widetilde{\mathbb{W}}^{1,p}_{0,\sigma}:=\widetilde{\mathbb{W}}_{0}^{1,p}(\mathscr{O})\cap\widetilde{\mathbb{L}}^{p}_{\sigma}$ and $\widetilde{\mathbb{W}}^{-1,p}_{\sigma}:=(\widetilde{\mathbb{W}}^{1,p^{\prime}}_{0,\sigma})^{\prime}$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$ for the dual space, where
$\widetilde{\mathbb{W}}^{1,p}_{0}(\mathscr{O}):=\begin{cases}\mathbb{W}^{1,p}_{0}(\mathscr{O})\cap\mathbb{W}^{1,2}_{0}(\mathscr{O}),&p\in[2,\infty)\\ \mathbb{W}^{1,p}_{0}(\mathscr{O})+\mathbb{W}^{1,2}_{0}(\mathscr{O}),&p\in(1,2).\end{cases}$ (2.7)
Lemma 2.2 ([52, Proposition 3.1]).
The Helmholtz projection $\mathscr{P}_{p}$ has a continuous linear extension $\widetilde{\mathscr{P}}_{p}:\widetilde{\mathbb{W}}^{-1,p}(\mathscr{O})\rightarrow\widetilde{\mathbb{W}}^{-1,p}_{\sigma}$ given by the restriction
$\widetilde{\mathscr{P}}_{p}\Phi=\Phi|_{\widetilde{\mathbb{W}}^{1,p^{\prime}}_{0,\sigma}},\;\;\text{ for }\;\;\Phi\in\widetilde{\mathbb{W}}^{-1,p}_{\sigma}\;\;\text{ with }\;\;\frac{1}{p}+\frac{1}{p^{\prime}}=1.$ (2.8)
Remark 2.2.
We will denote $\widetilde{\mathscr{P}}_{\frac{4}{3}}$ by $\mathscr{P}$ in the sequel.
@@ 2.3. Linear and nonlinear operators
$\mathscr{A}\boldsymbol{v}:=-\mathscr{P}\Delta\boldsymbol{v},\ \boldsymbol{v}\in\mathbb{X}.$
@@
[…]
Next, we define the trilinear form $b(\cdot,\cdot,\cdot):\mathbb{V}\times\mathbb{V}\times\mathbb{V}\rightarrow\mathbb{R}$ by
$b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{w})=\int_{\mathscr{O}}(\boldsymbol{u}(x)\cdot\nabla)\boldsymbol{v}(x)\cdot\boldsymbol{w}(x)\mathrm{d}x=\sum_{i,j=1}^{d}\int_{\mathscr{O}}\boldsymbol{u}_{i}(x)\frac{\partial\boldsymbol{v}_{j}(x)}{\partial x_{i}}\boldsymbol{w}_{j}(x)\mathrm{d}x.$
@@
$\left\{\begin{aligned} b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{v})&=0,\ \text{ for all }\ \boldsymbol{u},\boldsymbol{v}\in\mathbb{V},\\ b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{w})&=-b(\boldsymbol{u},\boldsymbol{w},\boldsymbol{v}),\ \text{ for all }\ \boldsymbol{u},\boldsymbol{v},\boldsymbol{w}\in\mathbb{V}.\end{aligned}\right$. (2.10)
The trilinear map $b:\mathbb{V}\times\mathbb{V}\times\mathbb{V}\rightarrow\mathbb{R}$ has a unique extension to a bounded trilinear map from $\mathbb{L}^{4}(\mathscr{O})\times\mathbb{W}^{1,4}(\mathscr{O})\times\mathbb{H}$ to $\mathbb{R}$. For any $\boldsymbol{u}\in\mathbb{W}^{1,4}(\mathscr{O})$ and $\boldsymbol{v}\in\mathbb{H}$, we have
@@
Let us now define the operator $\mathscr{J}(\boldsymbol{v}):=-\mathscr{P}[\mathrm{div}(\mathrm{A}(\boldsymbol{v})\mathrm{A}(\boldsymbol{v}))]$. Note that for $\boldsymbol{v}\in\mathbb{X}$, we have
$\|\mathscr{J}(\boldsymbol{v})\|_{\mathbb{X}^{\prime}}\leq C\|\mathrm{A}(\boldsymbol{v})\|^{2}_{4},$
and hence the map $\mathscr{J}(\cdot):\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ is well-defined.
Finally, we define the operator $\mathscr{K}(\boldsymbol{v}):=-\mathscr{P}[\mathrm{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))]$. It is immediate that
@@
Furthermore, the operator $\mathscr{K}(\cdot):\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ is well-defined.
@@
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$ and $\varepsilon_{0}:=1-\sqrt{\frac{\alpha^{2}}{2\beta\nu}}\in(0,1)$. Then, the operator $\mathscr{G}:\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ given by
@@
$\left<\mathscr{G}(\boldsymbol{y}_{1})-\mathscr{G}(\boldsymbol{y}_{2}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
@@
$\left<\nu\mathscr{A}\boldsymbol{y}_{1}-\nu\mathscr{A}\boldsymbol{y}_{2},\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>=\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}.$ (2.13)
@@
[…]
@@
$\left<\mathscr{B}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right$>
@@
From [38, Equation (2.21)], we have
[…]
@@
Now, from [38, Equation (2.13)], we have
@@
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$. Then, the operator $\mathscr{G}:\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ is demicontinuous.
@@
Let us take a sequence $\boldsymbol{y}_{m}\rightarrow\boldsymbol{y}$ in $\mathbb{X}$, that is, $\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{X}}=\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{V}}+\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\rightarrow 0$ as $m\rightarrow\infty$. For any $\boldsymbol{\phi}\in\mathbb{X}$, we consider
@@
$|\langle\mathscr{A}\boldsymbol{y}_{m}-\mathscr{A}\boldsymbol{y},\boldsymbol{\phi}\rangle|=|(\nabla(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi})|\leq\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{V}}\|\boldsymbol{\phi}\|_{\mathbb{V}}\rightarrow 0,\ \text{ as }\ m\rightarrow\infty,$ (2.16)
since $\boldsymbol{y}_{n}\rightarrow\boldsymbol{y}$ in $\mathbb{V}$. We estimate the term $\langle\mathscr{B}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.3) using (2.10), Hölder’s inequality and Lemma 2.1 as
@@
$\leq C\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}+C\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}+\boldsymbol{z}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}$
$\rightarrow 0,\ \text{ as }\ m\rightarrow\infty,$ (2.17)
since $\boldsymbol{y}_{m}\rightarrow\boldsymbol{y}$ in $\mathbb{X}$. We estimate the term $\langle\mathscr{J}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.3) using Hölder’s inequality as
$|\alpha\langle\mathscr{J}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle|$
@@
$\rightarrow 0,\ \text{ as }\ m\rightarrow\infty,$ (2.18)
since $\boldsymbol{y}_{m}\rightarrow\boldsymbol{y}$ in $\mathbb{W}^{1,4}(\mathscr{O})$. We estimate the term $\langle\mathscr{K}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.3) using Hölder’s inequality as
@@
$\rightarrow 0,\ \text{ as }\ m\rightarrow\infty,$ (2.19)
since $\boldsymbol{y}_{m}\rightarrow\boldsymbol{y}$ in $\mathbb{W}^{1,4}(\mathscr{O})$.
From the above convergences, it is immediate that $\langle\mathscr{G}(\boldsymbol{y}_{m})-\mathscr{G}(\boldsymbol{y}),\boldsymbol{\phi}\rangle\rightarrow 0$, for all $\boldsymbol{\phi}\in\mathbb{X}$. Hence the operator $\mathscr{G}:\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ is demicontinuous, which implies that, in addition, the operator $\mathscr{G}(\cdot)$ is hemicontinuous also. ∎
@@
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$. Then, the operator $\mathscr{G}:\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ is locally Lipschitz.